OpenStudy (anonymous):

What is the slope of a line parallel to 10x - 5y = 8? What is the slope of a line parallel to 4x + 2y = 5? What is the slope of a line perpendicular to x - 3y = 9? What is the slope of a line perpendicular to x - 5y = -10? What is the slope of a line parallel to 4x + y = -1? A) m = -5 B) m = 2 C) m = -3 D) m = -4 E) m = -2

OpenStudy (anonymous):

Are those the answers you suggest? And by slope does it mean gradient?

OpenStudy (anonymous):

matching find auto what is what

OpenStudy (anonymous):

*out

OpenStudy (nubeer):

hmm do you know how to identify the slope or gradient if the equation of line is given.. it's general form is y = mx +c... familiar with it?

OpenStudy (anonymous):

If you're trying to find the gradient of an equation, rearrange it so that it's in terms of y= something. For example, y=mx+c. Once it's in that form, the gradient is just whatever m is! For example, for the first one: \[10x-5y=8\] Add 5y to both sides, and take away 8 from both sides! \[5y=10x-8\] Divide both sides by 2! \[y=5x-4\] So the gradient is 5!

OpenStudy (nubeer):

aaahhh no.. don't you think it would be 2.. and why would you divide by 2? you have to make y isolate so you would divide by 5

OpenStudy (anonymous):

Oh sorry yeah, my bad! Divide both sides by 5 sorry, and you get y=2x-8/5, so the gradient is 2. Oops!

OpenStudy (anonymous):

10x - 5y = 8 slope = 2 (B) 4x + 2y = 5 slope = -2 (E) x - 3y = 9 perpendicular slope = -3 (C) x - 5y = -10 perpendicular slope = -5 (A) 4x + y = -1 slope = -4 (D)

OpenStudy (anonymous):

Perpendicular Slope = mPS Slope = m mPS = -1/m

OpenStudy (anonymous):

thats what i got too johana :D